OpenStudy (anonymous):

100x100÷2x100÷2x100÷21+34X321-332=? answering this will result in SUDDEN DEA medals :)

Nnesha (nnesha):

i'll give you medal if you answer it.

OpenStudy (anonymous):

yep but thats cheating a peice of paper and a pencil is better so NO CHEATING :c

OpenStudy (anonymous):

hint theres a remainder of nine :3

Nnesha (nnesha):

 multiply top fraction with the reciprocal of the bottom fraction example \[\huge\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b } \times \frac{ d }{ c }\]